OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence. Show all work 1, -4, -9, -14, . . .

OpenStudy (anonymous):

\[\sum_{0}^{11}(-5x+1)=-5*\sum_{0}^{11}x+\sum_{0}^{11}1=-5(0+1+2+3+...+11)+12*1=-5(11*12)/2+12=-330+12=-318\]

OpenStudy (anonymous):

Thank you SOOO much

OpenStudy (anonymous):

no problem :)